A rectangle has a length of 15
15
inches less than 5
5
times its width. If the area of the rectangle is 2750
2750
square inches, find the length of the rectangle.


Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

A rectangle has a length 15 inches less than 5 times its width. If the area of the rectangle is 2750 square inches, find the length.

Answer by ikleyn(52781)

.
A rectangle has a length 15 inches less than 5 times its width.
If the area of the rectangle is 2750 square inches, find the length.
~~~~~~~~~~~~~~~

Let w be the width; then the length is  (5w - 15) ft

The area equation is

    
    w*(5W-15) = 2750,

or

    w*(w-3) = 550  (after dividing both sides by 5).


You may solve this quadratic equation in a way you want,

or simply GUESS its solution mentally:  w = 25 ft.


So, the width is 25 ft;  the length is  5*25 - 15 = 110 ft.    ANSWER


CHECK.  25*110 = 2750 sq. feet,  the total area.   ! Correct !

Solved.


//////////////


In the future, do not allow, PLEASE, this wild way formatting in your posts . . .